OpenStudy (anonymous):

Find equation of line through (-2,-5) with slope of -1/7=m

OpenStudy (amistre64):

y --5 = -1/7.(x--2)

OpenStudy (amistre64):

y = (-1/7)x -(2/7) -5

OpenStudy (amistre64):

y = -x -37 ------- 7

OpenStudy (amistre64):

or -5 - (-1/7)(-2) = b -5 - 2/7 = b -37/7 = b y = -(1/7)x - 37/7

OpenStudy (amistre64):

or x + 7y = -37